OpenStudy (anonymous):

((x^2+5x+6)/3x )divided by (x+3)

OpenStudy (slaaibak):

(x+2)/3x

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

oh i ssee multiply by reciprical

OpenStudy (slaaibak):

(x^2 + 5x + 6)/3x = ((x+2)(x+3))/3x so the x+3 would cancel it out

OpenStudy (anonymous):

http://www.twiddla.com/548645

OpenStudy (amistre64):

x^2+5x+6 1 ---------- * --------- 3x (x+3) (x+3)(x+2) 1 ---------- * --------- 3x (x+3) x+2 ---- 3x

myininaya (myininaya):

\[\frac{x^2+5x+6}{3x} \div (x+3)=\frac{(x+2)(x+3)}{3x} \times \frac{1}{(x+3)}=\frac{x+2}{3x}\]

myininaya (myininaya):

omg where's the bike, amistre?